OpenStudy (anonymous):

A card is drawn at random from a standard deck of playing cards. You win $1 if you draw a seven and $24 if you draw an ace. Otherwise you loose $2. What is the expected value of the game?

OpenStudy (anonymous):

ishan u in which standard

OpenStudy (anonymous):

$22-$25 = -$3

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

52 cards

OpenStudy (anonymous):

ok have u given NTSE AND NSTSE

OpenStudy (anonymous):

no...whats that

OpenStudy (anonymous):

Probability of getting a 7 = 1/13 Probability of getting an Ace = 1/13 Probability of not winning = 11/13 Money totals: $2*1/13 + $24*1/13 -$2*11/13 = $4/13